Fermi National Laboratory
 

Operation Summary for October 15, 2010 - October 18, 2010

 

The Integrated Luminosity for the period from 10/11/10 to 10/18/10 was 53.26 pb-1.
NuMI reported receiving no protons on target during this same period due to target work.

Friday Oct 15
The day shift began with Operations monitoring the Tevatron (TeV) store 8174 (55.63E30) and the Recycler stash (325.8E10), with the Antiproton Source (Pbar) stacking (26.8E10), with Linac, Booster, and Main Injector (MI) operating normally, with the NuMI experiment off for target work, with MiniBooNE taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams.

During the day shift, Operations terminated store 8174 at 8:45 AM, put the TeV into shot setup, and established store 8175 at 10:10 AM with an average initial luminosity (AIL) of 294E30. MI and Pbar personnel conducted antiprotons TLG repetition rate studies.

During the evening shift, Operations established beam to the Meson FTBF CALICE experiment T-978 at 3:26 PM and disabled their beam at 10 PM. Operations terminated store 8175 at 10:43 PM, put the TeV into shot setup, and established store 8176 at 11:50 PM with an AIL of 285.66E30. MI and Pbar personnel completed their antiproton studies.

Saturday Oct 16
The midnight shift began with Operations monitoring store 8176 (245.91E30) and the stash (43.5E10), with Pbar stacking (17.1E10), and with MiniBooNE taking beam.

During the midnight shift, Operations reported some minor downtime due to various low-energy Linac RF stations tripping off. Operations established beam to T-978 at 6:30 AM.

During the day shift, Operations terminated store 8176 at 1:46 PM, put the TeV into shot setup, and established store 8177 at 3:20 PM with an AIL of 284E30. Operators rebooted a Pbar Accumulator PLC (A:PLC03) due to problems with some transfers to the Recycler. Operations reported some minor downtime when BRF15 tripped off on a modulator LCW flow fault; operators applied a mechanical transient to the flow meter to get the station running again.

During the evening shift, Operations disabled the T-978 beam at 6:07 PM.

Sunday Oct 17
The midnight shift began with Operations monitoring store 8177 and the stash, with Pbar stacking, and with MiniBooNE taking beam.

During the midnight shift, Operations terminated store 8177 at 3:48 AM, put the TeV into shot setup, and established store 8178 at 5:14 AM with an AIL of 284.7E30. Operations established beam to T-978 at 7:28 AM.

During the day shift, Operations monitored the machines.

During the evening shift, Operations terminated store 8178 at 6:26 PM, put the TeV into shot setup, and established store 8179 at 7:44 PM with an AIL of 293.4E30. Operations disabled the T-978 beam at 6 PM.

Monday Oct 18
The midnight shift began with Operations monitoring store 8179 and the stash, with Pbar stacking, and with MiniBooNE taking beam.

During the midnight shift, Operations established beam to T-978 at 6:30 AM. Operations reported some minor downtime due to various Linac and Booster trips.

The Plans for Monday
The plans for today are to stack, stash, and store.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 10/18/2010    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory